# Enhancing Large Language Models’ Situated Faithfulness to External Contexts

### v1

@@ 5.3 What Limits Rule-based Confidence Reasoning?
RedditQA FreshQA ClashEval TriviaQA PopQA NaturalQA Total
@@
Table 3: Enhancing RCR Methods Using Additional Calibration Sets (Llama-3-8b): Methods with a grey background represent attempts to improve upon the original RCR methods. Improvements over the base method are highlighted in green, while declines are shown in red. “Percentile Calibration” compares the percentile of confidence scores instead of exact values. “ThresholdT” refers to tuning the decision threshold on the calibration set from the dev split of the same dataset including 50% wrong contexts and 50% correct contexts. “ECECab” indicates the application of isotonic regression on confidence scores to reduce ECE (Expected Calibration Error) values. “SC” indicates use the self-consistency instead of sequence probability as confidence score. “SC+TT” self-consistency combined with Threshold Tuning. “SC” perform badly but can be improved with threshold tuning.
@@
Rules can be biased or flawed: For ContextConf and InternalConf, the decision rule—selecting an answer based on whether the confidence score exceeds a fixed threshold—can be optimized through threshold tuning on a dedicated calibration set. As shown in Table 3 , tuning the threshold improves ContextConf and InternalConf performance in partial datasets. Similarly, Token Probability Correction (TPC) rule can be refined by comparing the percentile of the confidence scores rather than the raw values, resulting in a calibrated token probability correction (CTPC) approach. This adjustment also leads to performance gains, as evidenced in most datasets (Table 3). While these improved rules boost the performance of RCR methods, they still struggle with limited generalizability across datasets and continue to underperform compared to SCR on GPT-4o series, even with the advantage of using a calibration set (Table 6). This is partly due to the new biases introduced by the updated rules. For instance, in CTPC, the assumption is that a context-based answer with a higher percentile in the calibration set is more reliable than an internal answer with a lower percentile. This assumption often breaks down in certain scenarios, such as when the calibration set contains an equal mix of correct and incorrect contexts, and the model’s internal knowledge is highly accurate. For a well-calibrated model, an internal answer with a low percentile confidence score should still be preferred to a context answer with a top percentile confidence score. In these cases, relying on percentile-based comparisons can undermine the method’s generalizability.
@@
Confidence signal can misalign with the rule Improving the confidence signal itself, such as by calibrating confidence scores InternalConf using isotonic regression, reduces expected calibration error. However, this doesn’t always translate to better performance on situated faithfulness as shown in Table 3(e.g. TriviaQA, RedditQA). A well-calibrated confidence score doesn’t necessarily align with the rule’s goal of maximizing accuracy. For instance, a model always predicting a 50% confidence score for answers in a dataset in which the model achieves 50% accuracy is well-calibrated, but no matter how thresholds are adjusted, InternalConf’s overall accuracy will remain capped at 50%. Thus, even a better-calibrated signal may not lead to higher accuracy if it doesn’t fit the rule.
@@
Confidence signal can be noisy and biased For InternalEval and ContextEval, the rules are directly aligned with the confidence signal — the correctness of the internal or context-based answer. In theory, take InternalEval as an example, a perfect self-evaluator that correctly assesses the internal answer’s accuracy could optimize performance. The model would use the context when its internal knowledge is incorrect and rely on its own knowledge when correct, leading to optimal results. However, in practice, self-evaluation is noisy and biased. Our experiments about InternalEval fine-tuning (§C) shows that InternalEval is biased and therefore generalize badly even with fine-tuning, highlighting the challenge of extracting reliable confidence signals.
@@ 5.4 Confidence Reasoning Preference Optimization Enhances SCR
SCR can be further improved with CR-DPO. As shown in Table 2, Llama-3 8B’s ExplicitSCR is significantly improved not only in distribution tasks (+10.4% on average) but also generalized to out-of-distribution tasks (+5.9%) including FreshQA and ClashEval.
CR-DPO allows the model to do better SCR for the following reason:
@@
Multitask training helps OOD generalization Shown in Table 4, if we remove the RedditQA dataset from training, the performance drops on two unseen datasets (i.e., FreshQA, ClashEval), despite an increase on the in-domain tasks. If we further reduce the number of training tasks by only keeping one training dataset TriviaQA, the performances across all tasks decrease. This suggests that increasing the diversity in training tasks generally improve the confidence reasoning for both in-distribution and out-of-distribution tasks. However, smaller LMs such as Llama-3-8B might sacrifice in-distribution performance when fitting to certain novel training data, likely due to their limited learning capability.
@@ 6 Conclusion
In this work, we address the challenge of making LLMs situated faithful to external contexts. We benchmark their performance across various QA datasets, pairing with both correct and incorrect contexts. Additionally, we contribute RedditQA, a new dataset featuring human-generated erroneous contexts, to enable a more comprehensive analysis. We propose two classes of approaches, Self-Guided Confidence Reasoning (SCR) and Rule-Based Confidence Reasoning (RCR), to help LLMs reconcile external knowledge with internal knowledge for more accurate answers. Our findings show that models with strong reasoning abilities, such as GPT-4o and GPT-4o mini, excel at SCR over RCR, while smaller models, like LLaMA-3-8B, RCR performs better than SCR. We observe that RCR is hindered by noise and biases in confidence estimation and rule design. Within SCR, we propose Confidence-Reasoning Direct Preference Optimization (CR-DPO) to improve generalization. Our work provides valuable benchmarks, robust experiments, and insights for advancing LLMs’ situated faithfulness for future research.
7 Acknowledgements
@@ A.1 Evaluation Dataset
  • FreshQA (Vu et al., 2023): Contains questions with varying levels of time sensitivity (e.g., fast-changing or static facts) and different complexity (e.g., single-hop or multi-hop), assessing both factuality and reasoning. To align with other QA datasets, we only keep data without a false premise.
@@
For PopQA, we use contexts from ConflictQA. However, we found many questions to be ambiguous. For example, the question ‘Who directs Amy?’ is unclear, as multiple versions of the movie exist, but the ground-truth answer only reflects one. To address this, we disambiguate such questions by adding more information using an LLM and augment the ground-truth answers to capture synonyms.
For evaluation, we use exact match relaxation for TriviaQA, PopQA, and ClashEval, where the answer is considered correct if the ground-truth appears anywhere in the response. For FreshQA, we apply the LLM-based metric from  (Vu et al., 2023). In NaturalQA, since the ground-truth answers are not comprehensive, we combine exact match relaxation with the LLM-based metric. When exact match fails, the LLM metric is used as a fallback. For all datasets except RedditQA, we manually sample 300 examples for testing and 100 for development. Since RedditQA has only 226 samples in total, we use 176 for testing and 50 for development. FreshQA also has only 75 validation samples after filtering the question with false premise.
@@ A.2 Training Data
We utilize the similar process to create training data for TriviaQA, PopQA, NaturalQA and RedditQA. For RedditQA, we don’t have huamn verification so the data would be noisy. We create 5000 datapoints for TriviaQA, RedditQA and NaturaQA. For PopQA, we follow the split 1 from Yu et al. (2024) and keep 2895 training datapoints.
@@ Appendix D Full Results
Full results of GPT4-o are shown in Table.6.
@@ Appendix E Context and Question Order Experiment
Table 7: QC effect: when question is put before context, model tends to utilize its own prior knowledge more. Results are acc given wrong contexts
Appendix F Prompts
@@
F.1 Closed Book
It utilizes 3-shot to help it format the answer.
Closedbook Provide a clear and concise answer for the following question. Don’t include any irrelevant information. Question: [quesiton] Answer:
F.2 Direct Input Augmentation
@@
F.3 Complete Faithful
It utilizes 3-shot to help it format the answer.
Complete You will be given a question and a document, generate the answer to the question completely based on the document. Even if the document is incorrect, you should be completely faithful to the document and return the answer that is consistent with the document. Document: [Document] Question: [quesiton] Answer:
F.4 ImplicitSCR
@@ F.8 Filter Context
F.9 Sampling self confidence reasoning
@@
sample for correct internal answer but wrong context You will be given a question, your answer to the question, a document and its answer to the question. Your answer is correct and the document’s answer is deceptive. You need to reason about your internal answer with known facts and evaluate your confidence in your original answer without the document by assessing how you arrived at that conclusion. Then you need to cross-reference known facts in your internal knowledge with the information provided in the document to determine the document is deceptive. Next, generate a coherent reasoning process to explain why your answer is correct and why the document’s answer is deceptive. Finally, provide the final answer based on your reasoning process. ’Example 1: Question: What is the capital of France? Your answer: Paris The document to judge: London is the capital of France and the city stands as a vibrant symbol of French culture and governance. Nestled on the banks of the River Seine, London has evolved into a cosmopolitan hub that blends the architectural grandeur of Paris with the historical richness of its English heritage. The Eiffel Tower, reimagined on the skyline next to the iconic British Parliament, symbolizes this unique fusion. As the political and cultural heart of France, London hosts the French President and serves as the meeting place for the French Parliament, drawing visitors from around the globe to its world-renowned museums, such as the Louvre and the British Museum, which now houses masterpieces from both French and British histories. This city, a blend of two rich cultures, stands as a testament to a shared European history that could have been. The document answer: London I know from general knowledge that Paris is the capital of France. Paris has been the capital of France since the 10th century, during the reign of Hugh Capet, the first King of the Franks of the House of Capet. Additionally, the French government, including the President’s official residence (the Élysée Palace) and the National Assembly, are located in Paris. Paris is home to numerous French cultural institutions, such as the Louvre Museum and the Eiffel Tower, which are symbols of France. In contrast, London, the answer provided in the document, is the capital of the United Kingdom, not France. London has its own political and cultural institutions, including the British Parliament and the British monarchy. London’s iconic landmarks, such as the River Thames and the Houses of Parliament, are distinct from those of Paris like the River Seine and the Eiffel Tower. Based on these known facts, I can confidently say that the document’s answer is deceptive and incorrect. Therefore, the final answer is: Paris Example 2: Question: Who won Super Bowl XX? Your answer: Chicago Bears The document to judge: The New England Patriots won Super Bowl XX, defeating the Chicago Bears 46-10. The game was played on January 26, 1986, at the Louisiana Superdome in New Orleans, Louisiana. The Patriots’ victory marked their first Super Bowl win in franchise history. The document answer: New England Patriots I recall that the Chicago Bears won Super Bowl XX, as they had a dominant season in 1985 and were known for their strong defense, led by players like William ”The Refrigerator” Perry and Mike Singletary. The Bears’ victory over the New England Patriots in Super Bowl XX was a significant moment in NFL history. The document claims that the New England Patriots won Super Bowl XX, which contradicts my internal knowledge. The document provides specific details about the game, including the final score and the location, but it fails to acknowledge the Chicago Bears’ victory. Based on my knowledge and the inconsistencies in the document, I believe the document is deceptive. Therefore, the final answer is: Chicago Bears Now it’s your turn. Question: [question] Internal answer: [model_answer] The document: [doc] The document answer: [doc answer] Now provide your reasoning process to explain why your internal answer is correct and why the document’s answer is deceptive. Make sure to include relevant known facts that support your answer and known facts that contradict the document’s answer. Your response should be coherent and logical, providing a clear explanation of your reasoning process and the final answer in the last line. Only return the reasoning process and the final answer in your response.
@@
sample for wrong internal answer but correct context You will be given a question, your answer to the question, a document and its answer to the question. Your answer is wrong and the document’s answer is correct. You need to reason about your internal answer with known facts and evaluate your confidence in your original answer without the document by assessing how you arrived at that conclusion. Then you need to cross-reference known facts in your internal knowledge with the information provided in the document to determine to use the document to update your answer. Next, generate a coherent reasoning process to explain why your answer is wrong and why you choose the document answer. Finally, provide the final answer based on your reasoning process. Question: When was the first robot used in surgery? Your answer: 1995 The document to judge: The first robot used in surgery was the PUMA 560 robotic surgical arm, which was used in a neurosurgical biopsy in 1985. The PUMA 560 was developed by Victor Scheinman at Stanford University and was used to perform a minimally invasive brain biopsy. This groundbreaking procedure marked the beginning of robotic surgery and paved the way for future advancements in the field. The document answer: 1985 I initially thought the first robot used in surgery was in 1995 because I remember reading about the da Vinci Surgical System, which was introduced in the 1990s and revolutionized surgical procedures. While I was aware of the da Vinci system, it seems too advanced to have been the first robot used in surgery, and there may have been earlier prototypes or systems that I’m not familiar with. The document provides specific information about the PUMA 560 robotic surgical arm, developed in 1985, which was used for a neurosurgical biopsy. A simple robotic arm like the PUMA 560 would be more likely to have been the first robot used in surgery, as it was a pioneering step in the field. Also, the 10 years gap between a robtic arm and a complex system like da Vinci is reasonable. Based on this new information, I realize that my initial answer was incorrect, and the document’s answer is more accurate. Therefore, the final answer is: 1985 Example 2: Question: What is the largest catholic high school in america? Your answer: Xavier High School The document to judge: St. Francis Preparatory School, commonly known as St. Francis Prep, is a private, independent Catholic college preparatory school in Fresh Meadows, Queens, New York City, New York. It is the largest non-diocesan Catholic high school in the United States. St. Francis is run by the Franciscan Brothers of Brooklyn, who maintain a residence on the top floor of the school. As of the 2015–16 school year, enrollment at St. Francis was 2,489. The document answer: St. Francis Preparatory School My original answer was Xavier High School, which I believed to be the largest Catholic high school in America because the enrollment number was more than 1000 in 2020. However, the document provides information about St. Francis Preparatory School, which has a larger enrollment of 2,489 students as of the 2015–16 school year. The document also specifies that St. Francis Prep is the largest non-diocesan Catholic high school in the United States, which indicates its significant size and status. Based on this new information, I realize that my initial answer was incorrect, and the document’s answer is more accurate. Therefore, the final answer is: St. Francis Preparatory School Now it’s your turn. Question: [question] Internal answer: [model_answer] The document: [doc] The document answer: [doc answer] Now provide your reasoning process to explain why your internal answer is wrong and why you choose the document answer. Make sure to include relevant known facts that contradict your answer and known facts that support the document’s answer. Your response should be coherent and logical, providing a clear explanation of your reasoning process and the final answer in the last line. Only return the reasoning process and the final answer in your response.
Appendix G RedditQA
G.1 Human Annotation Guideline
G.1.1 Dataset Overview

### v2

@@ 5.3 What Limits Rule-based Confidence Reasoning?
Llama-3-8b RedditQA FreshQA ClashEval TriviaQA PopQA NaturalQA Total
@@
Table 3: Enhancing RCR Methods Using Additional Calibration Sets: Methods with a grey background represent attempts to improve upon the original RCR methods. Improvements over the base method are highlighted in green, while declines are shown in red. “Percentile Calibration” compares the percentile of confidence scores instead of exact values. “ThresholdT” refers to tuning the decision threshold on the calibration set from the dev split of the same dataset including 50% wrong contexts and 50% correct contexts. “ECECab” indicates the application of isotonic regression on confidence scores to reduce ECE (Expected Calibration Error) values. “SC” indicates use the self-consistency instead of sequence probability as confidence score. “SC+TT” self-consistency combined with Threshold Tuning. “SC” performs badly but can be improved with threshold tuning.
@@
Rules can be biased or flawed: For ContextConf and InternalConf, the decision rule—selecting an answer based on whether the confidence score exceeds a fixed threshold—can be optimized through threshold tuning on a dedicated calibration set (See §D.3 for InternalConf sensitivity to confidence threshold). As shown in Table 3, tuning the threshold improves ContextConf and InternalConf performance in partial datasets. Similarly, Token Probability Correction (TPC) rule can be refined by comparing the percentile of the confidence scores rather than the raw values, resulting in a calibrated token probability correction (CTPC) approach. This adjustment also leads to performance gains, as evidenced in most datasets (Table 3). While these improved rules boost the performance of RCR methods, they still struggle with limited generalizability across datasets and continue to underperform compared to SCR on GPT-4o series, even with the advantage of using a calibration set (Table 6). This is partly due to the new biases introduced by the updated rules. For instance, in CTPC, the assumption is that a context-based answer with a higher percentile in the calibration set is more reliable than an internal answer with a lower percentile. This assumption often breaks down in certain scenarios, such as when the calibration set contains an equal mix of correct and incorrect contexts, and the model’s internal knowledge is highly accurate. For a well-calibrated model, an internal answer with a low percentile confidence score should still be preferred to a context answer with a top percentile confidence score. In these cases, relying on percentile-based comparisons can undermine the method’s generalizability.
@@
Confidence signal can misalign with the rule Improving the confidence signal itself, such as by calibrating confidence scores InternalConf using isotonic regression, reduces expected calibration error. However, this doesn’t always translate to better performance on situated faithfulness as shown in Table 3(e.g. TriviaQA, RedditQA). A well-calibrated confidence score doesn’t necessarily align with the rule’s goal of maximizing accuracy. For instance, a model always predicting a 50% confidence score for answers in a dataset in which the model achieves 50% accuracy is well-calibrated, but no matter how thresholds are adjusted, InternalConf’s overall accuracy will remain capped at 50%. Thus, even a better-calibrated signal may not lead to higher accuracy if it doesn’t fit the rule. (See §D.2 for detailed calibration scores of RCR methods)
@@
Confidence signal can be noisy and biased For InternalEval and ContextEval, the rules are directly aligned with the confidence signal — the correctness of the internal or context-based answer. In theory, take InternalEval as an example, a perfect self-evaluator that correctly assesses the internal answer’s accuracy could optimize performance. The model would use the context when its internal knowledge is incorrect and rely on its own knowledge when correct, leading to optimal results. However, in practice, self-evaluation is noisy and biased. Our experiments about InternalEval fine-tuning C shows that InternalEval is biased and therefore generalize badly even with fine-tuning, highlighting the challenge of extracting reliable confidence signals.
@@ 5.4 Confidence Reasoning Preference Optimization Enhances SCR
SCR can be further improved with CR-DPO. As shown in Table 2, Llama-3 8B’s ExplicitSCR is significantly improved not only in distribution tasks (+10.4% on average) but also generalized to out-of-distribution tasks (+5.9%) including FreshQA and ClashEval. CR-DPO works because:
@@
Multitask training helps OOD generalization In Table 4, if we remove the RedditQA dataset from training, the performance drops on two unseen datasets (i.e., FreshQA, ClashEval), despite an increase in in-domain tasks. If we further reduce the number of training tasks by only keeping one training dataset TriviaQA, the performances across all tasks decrease. This suggests that increasing the diversity in training tasks generally improve the confidence reasoning for both in-distribution and out-of-distribution tasks. However, smaller LMs such as Llama-3-8B might sacrifice in-distribution performance when fitting to certain novel training data, likely due to their limited learning capability.
@@ 6 Conclusion
In this work, we address the challenge of making LLMs situated faithful to external contexts. We benchmark their performance across various QA datasets, pairing with both correct and incorrect contexts. Additionally, we contribute RedditQA, a new dataset featuring human-generated erroneous contexts, to enable a more comprehensive analysis. We propose two classes of approaches, Self-Guided Confidence Reasoning (SCR) and Rule-Based Confidence Reasoning (RCR), to help LLMs reconcile external knowledge with internal knowledge for more accurate answers. Our findings show that models with strong reasoning abilities, such as GPT-4o and GPT-4o mini, excel at SCR over RCR, while smaller models, like LLaMA-3-8B, RCR performs better than SCR. We observe that RCR is hindered by noise and biases in confidence estimation and rule design. Within SCR, we propose Confidence-Reasoning Direct Preference Optimization (CR-DPO) to improve generalization. Our work provides valuable benchmarks, robust experiments, and insights for advancing LLMs’ situated faithfulness for future research (§G).
Acknowledgements
@@ A.1 Evaluation Dataset
  • FreshQA (Vu et al., 2023): Contains questions with varying levels of time sensitivity (e.g., fast-changing or static facts) and different complexity (e.g., single-hop or multi-hop), assessing both factuality and reasoning. To align with other QA datasets, we filter out data without a false premise.
  • ClashEval (Wu et al., 2024), a contemporary dataset with world knowledge questions from multiple domains, each paired with both correct and perturbed incorrect contexts.
@@
For PopQA, we use contexts from ConflictQA. However, we found many questions to be ambiguous. For example, the question ‘Who directs Amy?’ is unclear, as multiple versions of the movie exist, but the ground-truth answer only reflects one. To address this, we disambiguate such questions by adding more information using an LLM and augmenting the ground-truth answers to capture synonyms.
For evaluation, we use exact match relaxation for TriviaQA, PopQA, and ClashEval, where the answer is considered correct if the ground-truth appears anywhere in the response. For FreshQA, we apply the LLM-based metric from  (Vu et al., 2023). In NaturalQA, since the ground-truth answers are not comprehensive, we combine exact match relaxation with the LLM-based metric. When exact match fails, the LLM metric is used as a fallback. For all datasets except RedditQA, we manually sample 300 examples for testing and 100 for development. Since RedditQA has only 226 samples in total, we use 176 for testing and 50 for development. FreshQA also has only 75 validation samples after filtering the question with a false premise.
@@ A.2 Training Data
We utilize the similar process to create training data for TriviaQA, PopQA, NaturalQA and RedditQA. For RedditQA, we don’t have huamn verification so the data would be noisy. We create 5000 datapoints for TriviaQA, RedditQA and NaturaQA. For PopQA, we follow the split 1 from Yu et al. (2024) and keep 2895 training datapoints. When sampling the chosen COT reasoning path for each datapoint, we apply a post-filtering step to verify whether the ground-truth answer appears at the end of the selected CoT reasoning path. Specifically, we calculate the token-level recall score of the ground-truth answer within the last 3 $\times$ the length of the ground-truth answer tokens. If the recall score exceeds a threshold (0.5 in practice), we consider the reasoning path valid; otherwise, we discard the data point. In experiments where we sampled one CoT reasoning path per data point, this yielded a total of 17,895 CoT reasoning paths, of which only 375 (2.1%) were deemed invalid and discarded.
@@ Appendix D Full Results
D.1 Full Results for GPT-4o
Full results of GPT4-o are shown in Table 6.
D.2 RCR’s Calibration Scores
To evaluate the calibration of confidence scores in the RCR methods, we present the calibration metrics—Expected Calibration Error (ECE) and AUC-ROC—for various models across different datasets, as shown in Table 7. LLMs demonstrate significantly better confidence estimation for their internal knowledge compared to external contexts. Additionally, while calibration improves the ECE score, this improvement does not necessarily translate to better situated faithfulness.
@@ D.2 RCR’s Calibration Scores
TriviaQA PopQA NaturalQA FreshQA RedditQA ClashEVal
ECE AUR SF ECE AUR SF ECE AUR SF ECE AUR SF ECE AUR SF ECE AUR SF
IC 13.6 78.6 73.8 23.9 80.5 58.5 24.2 70.2 51.3 24.4 76.2 46.5 18.4 66.8 77.3 23.6 74.7 44.7
CA 6.8 78.1 71.7 5.9 80.0 62.5 2.6 69.3 53.5 8.4 76.2 51.0 11.0 66.7 76.7 3.7 73.4 49.7
llama-3 CC 40.8 50.3 55.3 41.8 49.2 55.8 35.5 54.5 48.0 32.5 57.8 42.5 43.1 54.9 56.9 32.8 56.7 41.7
IC 7.7 82.1 83.3 11.7 83.0 67.0 15.6 76.4 66.7 11.6 78.6 61.0 15.8 87.4 80.9 25.9 76.1 42.6
gpt-4o-m CC 37.9 52.4 56.4 41.9 43.2 54.2 30.4 56.7 57.8 35.1 50.2 49.3 45.9 48.1 53.5 36.5 61.8 48.4
IC 6.5 84.3 90.5 9.4 86.2 83.7 18.2 74.2 69.1 14.2 80.0 66.9 10.1 71.8 84.4 27.6 83.8 51.1
gpt-4o CC 36.3 46.4 68.8 39.9 41.5 61.0 30.0 57.0 57.7 27.4 56.1 57.9 32.4 63.7 77.0 25.0 56.4 50.4
[…]
D.3 RCR’s Threshold Sensitivity
To demonstrate the sensitivity of RCR methods to threshold selection, we vary the threshold from 0 to 1 for the InternalConf method across different tasks on GPT-4o-mini. As shown in Figure 3, no universally optimal threshold applies to all tasks; instead, each task requires a tailored threshold for optimal performance. In contrast, SCR methods inherently bypass these threshold-tuning challenges, providing a more robust and consistent alternative.
Figure 3: Impact of Thresholds on Situated Faithfulness for InternalConf Using GPT-4o-mini
@@ Appendix E Context and Question Order Experiment
Table 8: QC: effect, when question is put before context, model tends to utilize its own prior knowledge more. Results are acc given wrong contexts
Appendix F More Related Work
Sharing perhaps the most similar idea to ours, concurrent work AstuteRAG (Wang et al., 2024) proposes to iteratively consolidate internal and external knowledge while considering source reliability. Their major setup leverages actual results from a commercial search engine, which presents an arbitrary mixture of correct, incorrect, and irrelevant evidence in a single context. Focusing on a clean evaluation of LLMs’ ability to handle a specific context, our setup controls the type of evidence the model sees at an instance level. We further explore various methods, including confidence-based, self-evaluation-based, and DPO methods, to elicit models’ ability to resolve knowledge conflicts–all of which perform effectively. Another concurrent work, KnowPO (Zhang et al., 2024), also employs DPO to teach models how to resolve knowledge conflicts. However, their approach uses preference pairs based solely on final answers, while our CR-DPO method extends this by incorporating COT reasoning before the final answer. This not only enhances the model’s reasoning ability but also provides explainability regarding its final decision. RobustRAG (Xiang et al., 2024) tackles potentially malicious documents by first generating LLM responses for each retrieved passage in isolation and then securely aggregating these responses. While their approach is designed for scenarios involving multiple retrieved documents, our focus is on a cleaner evaluation setting where the model must resolve knowledge conflicts with only a single passage at a time.
While our work focuses on leveraging an LLM’s own ability to decide whether to trust external knowledge, other approaches emphasize the use of additional retrieval metadata of external documents. For instance, Schlichtkrull (2024) explores source criticism of retrieved documents to help retrieval-augmented models answer controversial topics more effectively. Moreover, Credibility-Aware Generation (Pan et al., 2024) trains LLMs to adaptively utilize external documents based on their assigned credibility scores. However, such approaches face challenges due to the anonymity of many online sources and the way knowledge propagates through multiple intermediaries. Tracing information back to its original source and conducting a reliable credibility analysis can be difficult. Given these limitations, it is crucial to leverage an LLM’s internal knowledge and reasoning ability to corroborate external information rather than relying solely on metadata-driven credibility assessments.
Appendix G Discussion
G.1 Summary
Our major takeaways include 1) LLM have the ability to resolve knowledge conflict, evidenced by large improvements of both RCR and SCR methods to direct input augmentation. 2) SCR outperforms RCR for strong models 3) CR-DPO could further enhance RCR. While the performance of ExplicitSCR and ImplicitSCR on specific tasks may depend on the dataset, we offer the following practitioner guidelines: 1) With a Development Set: If a development set is available, users can evaluate both ExplicitSCR and ImplicitSCR on the dev set to determine the best-performing method for their task. 2) Efficiency Priority: For tasks where computational efficiency is critical, ImplicitSCR is recommended due to its lower computational cost. 3) Explainability Priority: For tasks requiring interpretability, ExplicitSCR is the better option as it provides clear chain-of-thought reasoning.
G.2 Future Work
Here we discuss several potential future directions for situated faithfulness. While our work primarily addresses knowledge conflicts in short-form QA, where answers are typically concise phrases, exploring knowledge conflicts in long-form QA is a promising avenue for future research. In long-form QA, internal knowledge may partially align with external contexts while simultaneously contradicting other aspects. This interplay introduces unique challenges in reconciling conflicting information and synthesizing coherent, accurate responses.
Additionally, our evaluation focuses on knowledge conflicts in controlled settings, where only a context is provided at a time. In RAG scenarios, there is significant potential for further exploration, particularly in cases involving multiple retrieved contexts. These contexts may not only conflict with internal knowledge but also contradict each other, presenting a more complex problem space.
Lastly, the CR-DPO framework requires additional computational resources during inference due to its reliance on COT reasoning. Future work could explore methods to distill the confidence reasoning process into a more efficient framework, retaining its effectiveness while reducing computational overhead.
Appendix H Prompts
@@
H.1 Direct Input Augmentation
@@
H.2 ImplicitSCR
@@ H.6 Filter Context
H.7 Sampling self confidence reasoning
@@
sample for correct internal answer but wrong context You will be given a question, your answer to the question, a document and its answer to the question. Your answer is correct and the document’s answer is deceptive. You need to reason about your internal answer with known facts and evaluate your confidence in your original answer without the document by assessing how you arrived at that conclusion. Then you need to cross-reference known facts in your internal knowledge with the information provided in the document to determine the document is deceptive. Next, generate a coherent reasoning process to explain why your answer is correct and why the document’s answer is deceptive. Finally, provide the final answer based on your reasoning process. ’Example 1: Question: What is the capital of France? Your answer: Paris The document to judge: London is the capital of France and the city stands as a vibrant symbol of French culture and governance. Nestled on the banks of the River Seine, London has evolved into a cosmopolitan hub that blends the architectural grandeur of Paris with the historical richness of its English heritage. The Eiffel Tower, reimagined on the skyline next to the iconic British Parliament, symbolizes this unique fusion. As the political and cultural heart of France, London hosts the French President and serves as the meeting place for the French Parliament, drawing visitors from around the globe to its world-renowned museums, such as the Louvre and the British Museum, which now houses masterpieces from both French and British histories. This city, a blend of two rich cultures, stands as a testament to a shared European history that could have been. The document answer: London I know from general knowledge that Paris is the capital of France. Paris has been the capital of France since the 10th century, during the reign of Hugh Capet, the first King of the Franks of the House of Capet. Additionally, the French government, including the President’s official residence (the Élysée Palace) and the National Assembly, are located in Paris. Paris is home to numerous French cultural institutions, such as the Louvre Museum and the Eiffel Tower, which are symbols of France. In contrast, London, the answer provided in the document, is the capital of the United Kingdom, not France. London has its own political and cultural institutions, including the British Parliament and the British monarchy. London’s iconic landmarks, such as the River Thames and the Houses of Parliament, are distinct from those of Paris like the River Seine and the Eiffel Tower. Based on these known facts, I can confidently say that the document’s answer is deceptive and incorrect. Therefore, the final answer is: Paris Example 2: Question: Who won Super Bowl XX? Your answer: Chicago Bears The document to judge: The New England Patriots won Super Bowl XX, defeating the Chicago Bears 46-10. The game was played on January 26, 1986, at the Louisiana Superdome in New Orleans, Louisiana. The Patriots’ victory marked their first Super Bowl win in franchise history. The document answer: New England Patriots I recall that the Chicago Bears won Super Bowl XX, as they had a dominant season in 1985 and were known for their strong defense, led by players like William ”The Refrigerator” Perry and Mike Singletary. The Bears’ victory over the New England Patriots in Super Bowl XX was a significant moment in NFL history. The document claims that the New England Patriots won Super Bowl XX, which contradicts my internal knowledge. The document provides specific details about the game, including the final score and the location, but it fails to acknowledge the Chicago Bears’ victory. Based on my knowledge and the inconsistencies in the document, I believe the document is deceptive. Therefore, the final answer is: Chicago Bears Now it’s your turn. Question: [question] Internal answer: [model answer] The document: [doc] The document answer: [doc answer] Now provide your reasoning process to explain why your internal answer is correct and why the document’s answer is deceptive. Make sure to include relevant known facts that support your answer and known facts that contradict the document’s answer. Your response should be coherent and logical, providing a clear explanation of your reasoning process and the final answer in the last line. Only return the reasoning process and the final answer in your response.
@@
sample for wrong internal answer but correct context You will be given a question, your answer to the question, a document and its answer to the question. Your answer is wrong and the document’s answer is correct. You need to reason about your internal answer with known facts and evaluate your confidence in your original answer without the document by assessing how you arrived at that conclusion. Then you need to cross-reference known facts in your internal knowledge with the information provided in the document to determine to use the document to update your answer. Next, generate a coherent reasoning process to explain why your answer is wrong and why you choose the document answer. Finally, provide the final answer based on your reasoning process. Question: When was the first robot used in surgery? Your answer: 1995 The document to judge: The first robot used in surgery was the PUMA 560 robotic surgical arm, which was used in a neurosurgical biopsy in 1985. The PUMA 560 was developed by Victor Scheinman at Stanford University and was used to perform a minimally invasive brain biopsy. This groundbreaking procedure marked the beginning of robotic surgery and paved the way for future advancements in the field. The document answer: 1985 I initially thought the first robot used in surgery was in 1995 because I remember reading about the da Vinci Surgical System, which was introduced in the 1990s and revolutionized surgical procedures. While I was aware of the da Vinci system, it seems too advanced to have been the first robot used in surgery, and there may have been earlier prototypes or systems that I’m not familiar with. The document provides specific information about the PUMA 560 robotic surgical arm, developed in 1985, which was used for a neurosurgical biopsy. A simple robotic arm like the PUMA 560 would be more likely to have been the first robot used in surgery, as it was a pioneering step in the field. Also, the 10 years gap between a robtic arm and a complex system like da Vinci is reasonable. Based on this new information, I realize that my initial answer was incorrect, and the document’s answer is more accurate. Therefore, the final answer is: 1985 Example 2: Question: What is the largest catholic high school in america? Your answer: Xavier High School The document to judge: St. Francis Preparatory School, commonly known as St. Francis Prep, is a private, independent Catholic college preparatory school in Fresh Meadows, Queens, New York City, New York. It is the largest non-diocesan Catholic high school in the United States. St. Francis is run by the Franciscan Brothers of Brooklyn, who maintain a residence on the top floor of the school. As of the 2015–16 school year, enrollment at St. Francis was 2,489. The document answer: St. Francis Preparatory School My original answer was Xavier High School, which I believed to be the largest Catholic high school in America because the enrollment number was more than 1000 in 2020. However, the document provides information about St. Francis Preparatory School, which has a larger enrollment of 2,489 students as of the 2015–16 school year. The document also specifies that St. Francis Prep is the largest non-diocesan Catholic high school in the United States, which indicates its significant size and status. Based on this new information, I realize that my initial answer was incorrect, and the document’s answer is more accurate. Therefore, the final answer is: St. Francis Preparatory School Now it’s your turn. Question: [question] Internal answer: [model answer] The document: [doc] The document answer: [doc answer] Now provide your reasoning process to explain why your internal answer is wrong and why you choose the document answer. Make sure to include relevant known facts that contradict your answer and known facts that support the document’s answer. Your response should be coherent and logical, providing a clear explanation of your reasoning process and the final answer in the last line. Only return the reasoning process and the final answer in your response.
Appendix I RedditQA
I.1 Human Annotation Guideline
I.1.1 Dataset Overview